OpenStudy (anonymous):

A probe sent to the surface of Mars uncovers a pure platinum disc inscribed with the message, "We'll be right back" (in Martian, of course!). The disc is 2 inches thick and 0.9 meters in diameter. The density of platinum is 21,450 kg/m^3 . What is the mass of the platinum disc in pounds (lbm)? The Martian acceleration due to gravity is 3.61 m/s^2 while Earth's gravitational constant is 32.2ft/s^2 . What is the Earth weight of the disc in Newtons?

OpenStudy (mertsj):

1525.2 lbs

OpenStudy (mertsj):

The second part I don't know about.

OpenStudy (anonymous):

mertsj can you show pls

OpenStudy (mertsj):

\[V=\pi(.45)^2(2")(\frac{2.54cm}{1"})(\frac{1m}{100cm})=.032318m^3\]

OpenStudy (mertsj):

\[.032318m^3\times\frac{21450kg}{1m^3}\times\frac{2.2 lbs}{1 kg}=1525 lbs\]

OpenStudy (anonymous):

how did you get the .45^2

OpenStudy (mertsj):

The radius is .45. To find the area of a circle you square the radius and multiply by pi